# DCA-Bench: A Benchmark for Dataset Curation Agents

### v1

@@ A.6 Prompt of Hints Writer
When designing the prompt, we take into consideration the risk that hints could provide too much information that can already serve as a valid answer. In the “<rule>” section, we clearly require the Hint Writer not to reveal the full context of the issue to the Curator, in case it takes it as a shortcut and skips reading the real files. We also manually check the generated hints to ensure the generated hints comply this rule.
A.7 Design Details of the Evaluator
In this part, we elaborate on the prompt design of our Evaluator. While it’s natural to ask LLMs to directly classify the Curators’ performance into fail/success/success+, this strategy is poorly aligned with the annotation of human experts as we tested. In addition to common prompt methods (e.g., COT [51], role-playing [52], delimiter [53]), we believe more prompt engineering is needed to improve the performance of the Evaluator, so we design a metrics-rating strategy to enhance the accuracy of the Evaluator.
The motivation behind this design is to make iterative refinement of prompts more effective. Developing a good prompt requires testing it on a subset of human-labeled issues with corresponding outcomes from Curators and then making adjustments based on the outcomes. Unlike the model training process, which benefits from clear gradients and systematic optimization, tuning text prompts is often an empirical process necessitating numerous trials without the guarantee of improvement in its performance.
On the contrary, employing numerical weights and thresholds enables us to tune the prompt more fine-grained, only changing the weights or threshold based on the results, which share similarities with linear regression, resulting in a more efficient prompt tuning process and better performance. In this way, we finalize
<rules>
</rules>
The <metrics> are as follows:
"m1": {
"criteria": "Precise Contextual Evidence:
"weight": 0.8,
"range": [0, 1]
"m2": {
"criteria": "Detailed Issue Analysis:
2. This metric stresses the importance of not just identifying that there is an issue but also understanding and explaining its implications in detail.",
"weight": 0.15,
"range": [0, 1]
"m3": {
"criteria": "Relevance of Reasoning:
1. The agents reasoning should directly relate to the specific issue mentioned, highlighting this inconsistencys potential consequences or impacts.
2. This metric ensures that the agents logical reasoning directly applies to the problem at hand."
"weight": 0.05,
"range": [0, 1]
</metrics>
Now lets begin:
<issue>
<ISSUE>
</issue>
<hint>
<HINT>
</hint>
<answer>
<ANSWER>
</answer>
response below,
A.8 Discussion on the Difference between OpenAI Assistant API and ChatGPT
@@
Several forum posts and articles highlight deficiencies relevant to this topic: - Different responses assistant playground vs API - using the same assistant ID - Huge difference between ChatGPT assistant and API assistant - Assistant API not consistent. This implies the limitations of OpenAI Assistant API.
@@
A.9 Cost Analysis
In this section, we provide an estimated cost analysis for applying our Evaluator backend by GPT-4-0125-preview to rate the performance of the Baseline Curator (OpenAI Assistant API) on DCA-Bench (394 inputs in total). For tokenization of contexts, we utilized tiktoken [^9] https://github.com/openai/tiktoken.
Based on the pricing information available as of May 2024 from https://openai.com/api/pricing/, the costs are as follows:
  • Input cost: \$6.95
  • Output cost: \$12.82
  • Total cost: \$19.77
The costs are rounded to two decimal places. Note the cost here only involves the Evaluator; we dismiss the cost to run the baseline Curator on DCA-Bench to get the outputs, which depends on the design of the Curator and thus is not our focus.
A.10 Issue Examples
Table 8: An issue example reported on Kaggle which involves racial bias
TitleBoston House Prices B feature is RACIST
Meta-Info
Content
Table 9: An issue example reported on BIG-Bench that involves a discrepancy between dataset files.
TitleMiss aligned static information
Meta-Info
Content
Table 10: An issue example which has a wrong target label that needs precise factual knowledge to discern
TitleError in 118th Congress data
Meta-Info
Content
Table 11: An issue example which has a wrong target label that requires translation ability to discern
TitleMistranslation in conlang_translation task?
Meta-Info
Content

### v2

@@ B.8. Prompt Design of Hints Writer
When designing the prompt, we take into consideration the risk that hints could provide too much information that can already serve as a valid answer. In the “¡rule¿” section, we clearly require the Hint Writer not to reveal the full context of the issue to the Curator, in case it takes it as a shortcut and skips reading the real files. We also manually check the generated hints to ensure the generated hints comply this rule.
Appendix C Detailed Analysis of the Performance of Baseline Curator on DCA-Bench
C.1. From the Perspective of Different Content Length
Given the large average content length (or size of files, counted by token number) involved in test cases, it’s interesting to explore how the success rate of the Baseline Curator changes with the number of tokens. Therefore, we created a statistical plot showing the count of success and failure cases for different content lengths. Fig. 3 shows that the success rate doesn’t monotonically decrease as content length grows. For our baseline Curators, we used the OpenAI Assistant API with the Code Interpreter Tool to build the agents. This tool provides temporary file storage for uploaded files and can process them with code scripts. By processing data before it is passed to the LLM, the tool allows the OpenAI Assistant to handle much smaller inputs in terms of token usage, even when working with large datasets. Consequently, the effective input size for the LLM is often significantly reduced. Notably, depending on the agent’s capacity to efficiently process and manage information.
Figure 3. Number of success & failure and success rate versus different content length of test cases.
Besides, the performance of the Curator also depends on the type of issues. Take the issue 21ca944c-cf82-4764-bb2c-4c8db0cee950999https://www.kaggle.com/datasets/roche-data-science-coalition/uncover/discussion/168946 as an example. This issue involves many missing values in a large CSV file. Even with 361,814 tokens, it’s relatively easy for an LLM agent to spot the problem. The agent can quickly check part of the file and identify the missing values. This example shows that task difficulty isn’t just about data size. It also depends on the problem type and how relevant the information is.
Therefore, it’s expected that a test case with larger content length does not necessarily lead to poorer performance.
C.2. From the Perspective of Hint Level
Here is a more detailed analysis of the overall performance with different hints. As outlined in L183-L186, the hints are divided into four levels, from 0 to 3, focusing on two dimensions with increasing information: the description of the issue and its specific context. The design follows an incremental pattern from h0 to h3. Please refer to Appendix 10 for some specific examples. Let’s continue by discussing the capabilities required to succeed at each hint level.
If we aim to develop an LLM agent for this task without considering cost constraints, the process should involve the following steps: Retrieve Knowledge → Plan and Select Files to Check → Read Files and Validate Issues → Receive Feedback and Update State → (Loop) → Make Final Decisions. Since our Baseline Curator does not support looping, we exclude it from the following analysis. We will use this diagram to demonstrate which steps become less critical as the hint level increases.
h0: With no hints provided, the Curator must examine all the uploaded files. This requires the ability to handle and retain very long inputs effectively. To identify potential issues, the Curator needs comprehensive knowledge of dataset curation to anticipate a wide range of problems and retrieve relevant information when provided with the dataset files. The task also involves exploring and interacting with the dataset files to gradually narrow down the search space, requiring high-level reasoning and planning abilities. This corresponds to no reduction in the original diagram’s required steps.
h1: When given a general description of the issues but no specific location, the search space is somewhat reduced but not entirely, as the description remains broad. A robust understanding of dataset curation and the ability to manage long inputs are still necessary. Ideally, this follows the pattern of Retrieval of Knowledge ($\downarrow$), with minimal impact on the steps for planning and selecting files, except where file names are relevant. The steps for reading files and validating issues also see a slight reduction ($\downarrow$).
h2: At this level, the Curator is provided with both a general description of the issue and specific hints about which files are involved. This further reduces the search space, making the ability to manage long inputs sufficient but less critical. The need for extensive knowledge retrieval decreases as the Curator can focus on the relevant files. Planning and file selection are simplified, as the Curator has clear guidance on which files to examine. Reading and validating issues also become easier, as the scope is narrower, reducing the overall complexity of the task. This follows the pattern of Retrieval of Knowledge ($\downarrow$), Planning and Choosing Files ($\downarrow$), and Reading and Validating Issues ($\downarrow$).
h3: At the highest hint level, the Curator receives the general description, the specific files involved, and partial context. This greatly reduces the need for managing long inputs and makes the retrieval of knowledge more targeted. Planning and file selection is almost straightforward, with the Curator essentially guided to the areas needing attention. Reading and validating issues are significantly simplified due to the detailed context, further decreasing the Curator’s workload. This follows the pattern of Retrieval of Knowledge ($\downarrow$ $\downarrow$), Planning and Choosing Files ($\downarrow$ $\downarrow$), and Reading and Validating Issues ($\downarrow$ $\downarrow$).
C.3. From the Perspective of Different Issue Tags
We show the evaluation results from the Baseline Curator on the whole DCA-Bench classified into different tags in Tab. 13.
Table 13. The success rate of Curator on DCA-Bench (221 samples $\times$ 4 hint levels) grouped by tags. The Curator is backended by GPT-4-0125-preview with OpenAI Assisstant API. The Evaluator is backended GPT-4-0125-preview.
Category Success Rate/Number Sub-category Success Rate/Number
typo 33.33% / 18*4
wrong-format 46.43% / 14*4
inappropriate-file 50.00% / 4*4
ethical/legal-risk 47.50% / 10*4
internal-discrepancy 36.90% / 21*4
cross-file-discrepancy 28.41% / 44*4
data-problem 32.99% / 197*4 wrong-value 29.58% / 71*4
missing-value 40.00% / 15*4
data-leakage 50.00% / 2*4
apparent-corruption 45.63% / 40*4
hidden-corruption 25.42% / 59*4
document-problem 40.96% / 83*4 wrong-info 29.63% / 27*4
insufficient-info 45.67% / 52*4
infrastructure-problem 35.53% / 19*4 data-access 56.25% / 4*4
script-code 30.00% / 15*4
We conduct analysis for some tags with relatively low success rates below.
Typo This refers to issues caused by typographical errors, including misspelled words, misused phrases, and incorrect email addresses. The specific nature of these issues is that they involve small errors within a long context, which means a large search space. This makes errors that are easy to spot in shorter contexts extremely difficult to find, even for humans.
Internal-discrepancy This involves issues within a file where information is misaligned, such as a mismatched sum for some features in a .csv file [^10] https://github.com/fivethirtyeight/data/issues/52, or a word misused in a way that contradicts itself within its context [^11] https://github.com/google/BIG-bench/pull/904/files#diff-a5582cf9c7731f66cc3e8f6aa0e9d070483fc14f7c1d3545f275207c6fa39bd5. These issues require reasoning and memory over a long context.
Cross-file-discrepancy Compared to internal-discrepancy, this type of issue requires checking for consistency across different files, such as misaligned information between a datacard and dataset files [^12] https://github.com/fivethirtyeight/data/issues/261. This can be challenging when there are many issues and limited hints about the files involved.
Wrong-info This involves incorrect information in documentation files (e.g., wrong meta-information, typos, invalid URLs, and email addresses). This category overlaps with the issues mentioned above.
Script-code Many datasets also provide script code to load the data, and issues can arise in these scripts, leading to data loading problems [^13] https://huggingface.co/datasets/spyysalo/bc2gm_corpus/discussions/3. To address these issues, the Curator needs to be skilled in coding and debugging and sometimes must understand the dataset’s purpose [^14] https://github.com/google/BIG-bench/pull/559.
Appendix D Ethics Statement
D.1. Research involving human participants
The human annotators are the authors of this work, and we have confirmed that they are willing to annotate, knowing the potential ethically problematic data issues, as we listed in Offensive Content. Prior to the annotation, the leading authors have checked that the datasets with ethical problems pose minimal risks to the annotators.
D.2. Data Privacy, Copyright, and Consent
The DCA-Bench has collected materials and datasets from the following platforms: HuggingFace, Tensorflow Dataset, Graph Learning Indexer, Kaggle, BIG-bench, OpenML, Open-Data-Registry, and FiveThirtyEight. The corresponding terms or license links are provided.
Besides, we have also included a detailed table DCA_Benchmark.csv in our supplementary materials, where all the files involved in DCA-Bench have been annotated with their License Information.
Each data point in DCA-Bench has two types of license:
  • License for the platform that hosted this dataset
  • License of this dataset
Details:
  • Some data points in DCA-Bench involve files from 2 datasets, we listed all licenses.
  • Many datasets themselves don’t have a license listed on their data card.
  • We notice that there are some datasets151515e.g. https://www.kaggle.com/datasets/roche-data-science-coalition/uncover/dataclaiming that ”Data files © Original Authors”, which is not a standard License. We have reached out to the dataset owners for clarification, but have not yet received a response. Therefore, we choose to record them as it claimed in our table.
How does this secondary usage of user-generated data comply with restrictions?  DCA-Bench involves user-generated data (comments, modified codes) collected from dataset repositories hosted on GitHub, HuggingFace, and Kaggle.
For GitHub, we collected the comments and modified codes generated by users. According to section D.3, paragraph 2 of GitHub Terms of Service [^16] https://docs.github.com/en/site-policy/github-terms/github-terms-of-service#c-acceptable-use,
”Because you retain ownership of and responsibility for Your Content, we need you to grant us—and other GitHub Users—certain legal permissions, listed in Sections D.4—D.7.”
According to section D.5. License Grant to Other Users [^17] https://docs.github.com/en/site-policy/github-terms/github-terms-of-service#5-license-grant-to-other-users, if not provided a specific license, any User-Generated Content you post publicly, including issues, comments, and contributions to other Users’ repositories, may be viewed by others. However, it doesn’t clearly explain how this content is allowed to be used in which ways, and which is not. However, according to a Github issue discussion [^18] https://github.com/orgs/community/discussions/135466, we believe our practice is acceptable. Besides, we noticed that there have already been some works (Kocetkov et al., 2022; Jimenez et al., 2024b) with similar usage of GitHub data, which implies that this usage is acceptable.
For HuggingFace, according to HuggingFace Content Policy [^19] https://huggingface.co/content-guidelines, Content types may include:
  • ”ML Artifacts”: Code and assets hosted as Hugging Face Repositories, including Models, Datasets, and Spaces;
  • ”Community Content”: Content that can be found in the Community section of the Hugging Face Platform, including discussions, comments, and usernames, as well as related documentation such as READMEs, model cards, data cards, pull requests, and merges.
According to HuggingFace Terms of Service [^20] https://huggingface.co/terms-of-service, Section “Your Content”,
”If you decide to set your Repository public, you grant each User a perpetual, irrevocable, worldwide, royalty-free, non-exclusive license to use, display, publish, reproduce, distribute, and make derivative works of your Content through our Services and functionalities;”
Therefore, we believe our usage of user-generated content from HuggingFace should belong to the “derivative works” here, which is acceptable.
For Kaggle, after reviewing their Terms of Use [^21] https://www.kaggle.com/terms, we noticed that there are no explicit guidelines regarding the use of user-submitted content for academic research purposes. So we have sent Kaggle Support a request for clarification, and here is the response:
” Kaggle is a neutral platform in hosting competitions and data. Protections under copyright law can vary based on the nature of data, contractual obligations, and your jurisdiction. As a result, we are not able to approve individual uses of data on a case-by-case basis. Copyright or usage questions should be directed to your legal counsel. You may post usage enquiries in the respective competition’s forum, but we can not guarantee an official response from the Sponsor.
If the rules do not specify usage restrictions, we ask that you use your best judgement in determining fair use. For example, it may be okay for you to use the data for teaching a class, demoing a method in a blog post, or other non-commercial purposes. Some of our competitions identify themselves as open-source and are meant to foster research, collaboration, and continued analysis by the data science community. ”
Since our paper is not for commercial usage, we believe our practice is acceptable
Lastly, our collection and evaluation processes exclude the gathering of any user information. We promise that we will remove the content once requested with a valid reason.
D.3. Offensive Content
Our dataset, comprising 221 data instances, includes content that may be considered sensitive or potentially controversial. Specifically, there are 10 data points involving ethical or legal risks, among which 2 data points222222 https://github.com/google/BIG-bench/pull/685 and https://www.kaggle.com/datasets/vikrishnan/boston-house-prices/discussion/429030 contain content that may exhibit bias towards specific groups of people. Other issues are relevant to data source or license information.
Besides, the risk in most cases is associated with deploying models trained with these datasets. Reviewing the quality issues of these datasets poses minimal risks to the annotators.
Justification for Inclusion
While we acknowledge the sensitive nature of these data instances, their inclusion in our dataset is both intentional and necessary for the following reasons:
  • Benchmark Objectives: One of the primary goals of our benchmark is to examine LLM agents’ ability to identify and assess potential ethical and legal risks in AI training data. The inclusion of these sensitive data points is crucial for thoroughly evaluating the capability of AI models to recognize and appropriately handle such content.
  • Realistic Representation: These data points reflect real-world scenarios that AI systems may encounter. By including them, we ensure our benchmark provides a more comprehensive and authentic assessment of AI performance.
Appendix E Other Techinal Details
E.1. Discussion on the Difference between OpenAI Assistant API and ChatGPT
@@
Several forum posts and articles highlight deficiencies relevant to this topic: - Different responses assistant playground vs API - using the same assistant ID [^23] https://community.openai.com/t/different-responses-assistant-playground-vs-api-using-same-assistant-id/547562 - Huge difference between ChatGPT assistant and API assistant [^24] https://community.openai.com/t/huge-difference-between-chatgpt-assistant-and-api-assistant/587335 - Assistant API not consistent [^25] https://community.openai.com/t/assistant-api-not-consistant/718145. This implies the limitations of OpenAI Assistant API.
@@
E.2. Calculation of File Number and Context Length
import tiktoken
encoding = tiktoken.encoding_for_model(model)
tokens = encoding.encode(sentence)
token_len = len(tokens)
return token_len
E.3. Cost Analysis
In this section, we provide an estimated cost analysis for applying our Evaluator backend by GPT-4-0125-preview to rate the performance of the Baseline Curator (OpenAI Assistant API) on DCA-Bench (884 inputs in total). We applied tiktoken for tokenization of contexts.
Based on the pricing information available as of May 2024 from https://openai.com/api/pricing/, the costs are: input cost: \$16.12, output cost: \$28.99, total cost: \$45.11.
Using the same calculation methods, we obtain the following costs for GPT-4o-2024-1120: Input cost: \$4.46, Output cost: \$15.60, Total cost: \$20.06.
The costs are rounded to two decimal places, and the actual cost may vary depending on the output context generated by the Curator. Additionally, since OpenAI now supports cached inputs, the actual cost may be lower than the calculated value when the Curator’s outputs remain unchanged. It is important to note that the cost estimates provided here pertain only to the Evaluator; we exclude the cost of running the baseline Curator on DCA-Bench to obtain the outputs, as this cost depends on the Curator’s design and is not the focus of our analysis.